ganeshie8 (ganeshie8):

Show that \[\sum\limits_{n=1}^{\infty}\dfrac{(-1)^{n+1}}{n^s}=\frac{2^{s-1}-1}{2^{s-1}}\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}\] for all integers \(s>1\)

myininaya (myininaya):

eww do we have to use the reimaann zeta function thingy ?

myininaya (myininaya):

and however it is spelled

myininaya (myininaya):

hmmm maybe we can use induction

myininaya (myininaya):

that still might require that zeta function thing

ganeshie8 (ganeshie8):

Oh I think induction should work xD but i havent tried induction yet... it is a relation between riemann zeta and dirichlet eta functions but that relation can be derived using algebra w/o mentioning about zeta..

ganeshie8 (ganeshie8):

came across this nice relation while working on @xapproachesinfinity 's latest problem..

OpenStudy (zarkon):

expand the right hand side...works out pretty easily

OpenStudy (thadyoung):

@bibby

ganeshie8 (ganeshie8):

Wow yes that should be the simplest way to work this \[\begin{align}\frac{2^{s-1}-1}{2^{s-1}}\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}&=\left(1-\frac{1}{2^{s-1}}\right)\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}\\~\\ &=\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}-\frac{1}{2^{s-1}}\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}\\~\\ &=\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}-\sum\limits_{n=1}^{\infty}\dfrac{2}{(2n)^s}\\~\\ &=\left(\sum\limits_{n=1}^{\infty}\dfrac{1}{n^s}-\sum\limits_{n=1}^{\infty}\dfrac{1}{(2n)^s}\right) - \sum\limits_{n=1}^{\infty}\dfrac{1}{(2n)^s} \\~\\ &=\left(\sum\limits_{n=1}^{\infty}\dfrac{1}{(2n-1)^s}\right) - \sum\limits_{n=1}^{\infty}\dfrac{1}{(2n)^s} \\~\\ &=\sum\limits_{n=1}^{\infty}\dfrac{(-1)^{n+1}}{n^s} \\~\\ \end{align}\]